OpenStudy (anonymous):

This one seems hard lol. Question: Let f(x) = 5x + 12. Find f^−1(x). Answer Choices: A) 12x + 5 B) 1 OVER 5x + 12 (Fraction) C) 5 OVER x-12 (Fraction) D) x-12 OVER 5 (Fraction)

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

Its f raised to the negative 1x power, right?

OpenStudy (anonymous):

Yes.

OpenStudy (maddy1251):

Do you want to attempt it, and then see what you can come up with? I'm gonn try working it out :-)

OpenStudy (anonymous):

Hmm, I'm already confused at the start with how I would being the process lol.

OpenStudy (anonymous):

I brought more paper though to take notes. xD

OpenStudy (maddy1251):

At least you are prepared xD I had to go through my notes to figure this out. One second.

OpenStudy (anonymous):

Okay lol, take your time. :)

OpenStudy (maddy1251):

Okay, so, we want to find the inverse of the function to solve for this. Big words, seems hard, actually really easy.

OpenStudy (maddy1251):

Lets start by setting f(x)=y, so that way it will be easy.

OpenStudy (anonymous):

So we'd change the f(x) to y as in: y = 5x + 12 ??

OpenStudy (maddy1251):

y=5x+12. This is what it will look like. Yes.

OpenStudy (maddy1251):

Then, switch the x to a y in 5x+12, looking like x=5y+12.

OpenStudy (maddy1251):

Got that much?

OpenStudy (anonymous):

Yup, I understand it so far! :)

OpenStudy (maddy1251):

Okay, now we want to solve for y , like we did for the other problems. :-)

OpenStudy (maddy1251):

(f^-1)(x)=(x-12)/5

OpenStudy (anonymous):

How did you do that? Lol

OpenStudy (maddy1251):

x=5y+12 x-12=5y/5 (x-12)/5=x

OpenStudy (maddy1251):

We want to isolate the variable. I meant it equals y, not x. Force of habit. :P

OpenStudy (maddy1251):

Subtract 12 from both sides, that's the easiest way to start.

OpenStudy (maddy1251):

x=5y+12 -12 -12 Rule of math forever, what we do to one side, we do to the other. x-12=5y

OpenStudy (maddy1251):

However, y is not by itself yet. For it to be isolated, or by itself, we then divide by 5, doing the inverse of what 5 is doing to y. Soooo, like I said, what you do to one side, we do to another. (x-12)/5=5y/5

OpenStudy (maddy1251):

(x-12)/5=y, or f(x), since we changed y to f(x) for simple reasons.

OpenStudy (anonymous):

Oh okay so since y=5 right now (5y) let's say it were 8y. Would you be this way: (x-12)/8=8y/8. So you're basically taking the y and dividing on each side then?

OpenStudy (maddy1251):

You're dividing by the number value, not the variable. But, you typed it correctly so I am assuming you know how to do it :-), We want 'y' alone, so we divided by 5, not y.

OpenStudy (maddy1251):

If we divided by 5, we would of had (x-12)/y=5

OpenStudy (maddy1251):

I mean *y*

OpenStudy (anonymous):

Oh, xD Yes, that's what I meant, the 5 on each side, not the y. But I meant the number associated with the y variable. lol

OpenStudy (maddy1251):

Yep, because we were solving for 'y', and since subtracting 12 wasn't enough to get that little y variable alone, we divided by 5, because that was the opposite of what was being applied to y in this case. And you do it to each side, because that's that rule of math :)

OpenStudy (maddy1251):

5 had the attatchment to the y through multiplication, but 5 needed to go to the side where the 12 went, because y needed to be alone.

OpenStudy (maddy1251):

So, we divided by 5. inverse operations.

OpenStudy (anonymous):

I see, I understand now. I thought it was going to be really tricky lol. So, what would we do next? I feel like x-12/5 is the answer, but are there any steps left?

OpenStudy (maddy1251):

Nope! There is nothing to be done, because the question asks to find f(x), which we did through all that lovely math. That is your final answer, since no substitution is needed, such as plugging in a variable such as 2.

OpenStudy (anonymous):

Alright then! I'm glad we went over this one! I only have one left! :) I feel confident about all my answer because I am receiving help which makes me not so anxious haha. I'll tag you in the next one, once again, thank you. :)

OpenStudy (maddy1251):

Me too! Its good review for me, too :) Double winning. You got this.